1.A hospitalized patient with heart failure is on a 1000 mL fluid restriction each 24 hours. During the 7 a.m. to 7 p.m. shift he has three meals, consuming 1/2 cup of coffee with each. He also has 2 oz. of water each time he takes his medications at 7 a.m., noon, and 5 p.m. In the Midafternoon, he drinks 240 mL of soda. How many milliliters may he consume for the rest of the remaining hours? ______mL.

**(Read carefully)

2. Divide these fractions: 2/3 ÷ 3/4. Convert to the lowest terms. (Two thirds divided by three fourths.)

a. 6/12

b, 1/2

c. 5/12

d. 8/9

3.The correct conversion for the improper fraction to a mixed number for 77/9 is
Answer #1

1, Fluid restriction 1000ml
7 am to 7 pm three times meals with 1/2 cup of coffee with each=1/2cup X 3times
1/2cup =125mlX3 = 375ml

water 2oz with each medication three times
2oz=60mlX3=180ml
Midafternoon =240ml soda
total =375+180+240= 795ml
1000ml-795ml=205ml he may consume for the rest of the remainaing hours.
2, 2/3 divided by 3/4
2/3 =0.7
3/4=0.8
0.7 divided by0.8 = 0.875=0.8
0.8 =8/9
Ans = d
3, Improper fraction =77/9
Mixed fraction = whole number +proper fraction
77=8X9+5
(8X9+5)/9 = 8+5/9 = mixed number